OpenStudy (anonymous):

What operation must be done to both sides of the equation to solve p - -3 7/9 = 5/6 Add five-sixths Subtract three and seven-ninths Subtract five-sixths Add three and seven-ninths

OpenStudy (lncognlto):

Just to clarify the equation: p - (-3 and 7/9) = 5/6, right?

OpenStudy (anonymous):

@lncognlto no let me re write it..

OpenStudy (anonymous):

\[p- -3\frac{ 7 }{ 9 } = \frac{ 5 }{ 6 }\]

OpenStudy (lncognlto):

So then \[p + 3\frac{ 7 }{ 9 } = \frac{ 5 }{ 6 }\]

OpenStudy (lncognlto):

Then one would need to subtract 3 and seven ninths from each side of the equation.

OpenStudy (anonymous):

@lncognlto ty for helping!

OpenStudy (lncognlto):

yw :)